# Conceptualization, Methodology, Software, Analysis, Writing and reviewing manuscript

### v1

@@ 3 Numerical Implementation of Viscoelasticity via Energy-based PINN
This section details the numerical framework used to simulate the nonlinear viscoelastic phenomena described in the previous section. We employ an energy-based PINN approach, also known as the Deep Energy Method (DEM) [42, 23, 18, 17]. This method seeks the admissible displacement field that minimizes the total potential energy of the system. To handle the inherent time-dependency effect of viscoelasticity and growth, we utilize an incremental, time-stepping procedure, illustrated schematically in Fig. 3.
@@ 3.1 Energy-Based Formulation for Viscoelasticity
where $N(\cdot)$ is the neural network mapping with trainable parameters $\mathbf{\varphi}$ at time step $t_{i}$, and $\mathbf{x}$ is the position of material points. In practice, the neural network architecture, i.e. the number of hidden layers and the number of nodes per layer, is flexible and can be adjusted based on the PDE complexity [43, 44, 45, 46]. In this study, the neural network has three hidden layers and each layer contains twenty nodes.
@@
using the gradient descent method (e.g., Adam [47]). Once converged, the displacement field $\mathbf{u}(\mathbf{x},t_{i})$ is obtained, and the corresponding stress tensor can be calculated using the constitutive equation. Then, the simulation advances to the next time step $t_{i+1}$ with increment $\Delta t$. This incremental process allows simulation of the full time-dependent viscoelastic response.
@@ 3.2 Implementation of Boundary Conditions
Traction boundary conditions are naturally incorporated through the external work term in the potential energy functional $\Pi$. For displacement (essential) boundary conditions, there are two main approaches to impose within energy-based PINN: the soft and hard ways. The soft way means adding a penalty term to the loss function that penalizes deviations from the prescribed displacement [48, 17]. We use the hard enforcement to impose the displacement boundary conditions. This method imposes the boundary conditions exactly without adding penalty terms to the loss function, which can be beneficial for training convergence [17]. Using this method, the output is then modified as:
@@
where $\odot$ means the element-wise product, $b$ is an approximate distance function to the boundary $\Gamma(\mathbf{x})$ which denotes the shortest distance of a point $\mathbf{x}$ to the essential boundary, and $\mathbf{\bar{u}}(\mathbf{x},t)$ is a function that can satisfy the essential boundary conditions when $\mathbf{x}\in\Gamma(\mathbf{x})$ The distance function is non-negative and $b(\mathbf{x})=0$ only when $\mathbf{x}\in\Gamma(\mathbf{x})$. For the simple geometry with explicit boundaries, the method in Eq. (20) is easy to implement. For complex geometries, the boundary conditions can be imposed using distance functions [49]. Different methods to impose displacement boundary conditions on PINN have been listed and compared in [50].
@@ 3.3 Benchmark Test and Discussion
Figure 4: Viscoelastic creep and stress relaxation cases solved by PINN. (a) The schematic diagram of stress relaxation of a cantilever. The left boundary is fixed, and the right boundary is fixed horizontally after stretching for a displacement $\Delta L$. (b) Comparison of the stress relaxation results at the time $t=3\tau$ from FEM and PINN. (c) The schematic diagram of the creep of a cantilever. The left boundary is fixed, and the right boundary is subjected to a horizontal, uniformly distributed tensile force. (d) Comparison of the creep results at the time $t=3\tau$ from FEM and PINN. (e) The training dynamics of the energy-based PINN. The potential energy loss decreases as training progresses. Small oscillations may occur during the convergence process, which are brought in by the optimization algorithm itself.
@@
The viscoelastic creep displacement obtained from the PINN agrees well with the finite element results (Fig. 4d). These benchmark results demonstrate that the implemented energy-based PINN framework, using an incremental time-stepping approach, accurately captures the fundamental time-dependent responses of viscoelastic materials.
@@ 3.4 Oscillatory Convergence of Neural Network
An interesting characteristic observed during the training of energy-based PINN is the oscillatory behavior in the convergence of the loss function or predicted displacements (Fig. 4e). These oscillations arise from the dynamics of the gradient descent optimizer (e.g., SGD and Adam), which is very common during deep learning. Such oscillations can facilitate escape from poor local minima, but they may impede stable convergence during later optimization stages. In contrast, such oscillatory behavior may be advantageous for capturing phenomena like buckling in viscoelastic materials.
Classical numerical methods often require explicit artificial perturbations or initial imperfections to initiate buckling bifurcation from a perfect equilibrium path. We hypothesize that the inherent training oscillations within the energy-based PINN can act as a form of natural, internal perturbation. As the optimizer explores the energy landscape near a bifurcation point (where the landscape topology changes), these numerical oscillations can be sufficient to nudge the solution off the unstable primary path and guide it towards the physically stable, lower-energy buckled configuration. The potential applications are explored further in the next sections.
@@ 4.1 Creep Buckling of a Viscoelastic Cantilever Beam
Figure 5: Creep buckling of a cantilever beam under axial compression. (a) The schematic diagram of a compressed cantilever beam. (b) The uniform discretization of the beam. (c) The displacement in the y-direction of the creep buckling cantilever beam at different times obtained from PINN under pressure $2\times 10^{-2}$ Pa, with Adam optimizer and learning rate $5\times 10^{-3}$. (d) The deformed cantilever beam obtained from FEM under pressure $2\times 10^{-2}$ Pa, with initial geometric imperfections.
@@
The evolution of the deformation over time is shown in Fig. 5c. As creep deformation progresses, the beam undergoes significant lateral displacement, leading to buckling. The creep buckling with FEM is predicted by adding some initial geometric imperfections (Fig. 5d) under the same axial pressure. The PINN solution captures the gradual accumulation of strain and the onset of instability. To validate the predicted buckling behavior, we compare the eigenvalue of the viscoelastic cantilever with two reference cases obtained from FEM simulations (ABAQUS, plane strain analysis with CPE8R elements): one is modeled as an elastic material with initial modulus $E_{0}=E_{\infty}+E_{1}$, representing the upper bound of stiffness, and another is modeled with its long-term modulus $E=E_{\infty}$, serving as a baseline reference. The eigenvalue of a beam with the infinity modulus ($E_{\infty}$) is obtained from a finite element simulation, which is $9.36\times 10^{-3}$, and the beam with initial stiffness ($E_{0}$) is $2.34\times 10^{-2}$. The buckling pressure $2\times 10^{-2}$ for a viscoelastic cantilever is between these two elastic conditions. These predictions confirm that the buckling pressure of the viscoelastic cantilever falls within the expected range between the two elastic cases. Crucially, the buckling behavior in the PINN simulation emerged without introducing any explicit geometric imperfections or transient load perturbations. The transition to the buckled state appears to be triggered by the inherent dynamics of the optimizer exploring the solution space near the bifurcation point, consistent with the hypothesis outlined in Section 3.4. As supporting evidence, simulations performed with a sub-critical load ($p=5\times 10^{-3}<p_{\text{cr}}$) showed only minor oscillations or shaking during training without leading to buckling, as detailed in Appendix B. This highlights a potential advantage of PINN for stability analysis.
@@ 4.2 Creep Buckling of a Thin Cylindrical Shell under External Pressure
The discretization of the shell domain is shown in Fig. 6b, where the uniform sampling points are used to resolve the spatial dependence of the displacement field. Sample points on the outer surface of the shell are used to calculate the outer surface displacement. The implementation of this problem in PINN is under the cylindrical coordinate. The boundary condition is introduced to the neural network as
@@
Figure 6: Creep buckling of a thin cylinder with outside radial pressure. (a) The schematic diagram of a compressed quarter cantilever cylinder, with the pressure equal to $1\times 10^{-2}$. (b) The discretization of the cylinder and the integration at the outer surface. (c) The outer surfaces at different times. (c) Deformed quarter cylinder at different times. The forth order eigenvalue of a beam with the infinity modulus ($E_{\infty}$) is obtained from a finite element simulation (Plain strain, CPE8R), which is $6.44\times 10^{-3}$, and the cylindrical shell with initial stiffness ($E_{0}$) is $1.61\times 10^{-2}$. The buckling pressure $1\times 10^{-2}$ for a viscoelastic cantilever is between these two elastic conditions.
@@
The evolution of the shell’s deformation is presented in Figs. 6c and 6d. In Fig. 6c, the deformation profile of the outer surface is plotted at different time instances, illustrating the progressive lateral deflection and the onset of buckling. The initially circular shell gradually deviates from its original shape due to viscoelastic creep, eventually leading to instability. The corresponding displacement fields at specific time snapshots are visualized in Fig. 6d. At $t=0^{+}$, the shell exhibits a nearly uniform radial contraction due to the applied pressure. As creep deformation progresses, localized bending deformations emerge, which intensify over time, culminating in a pronounced buckled shape at $t=4\tau$. To validate the PINN predictions, the buckling pressure obtained from the viscoelastic simulation is still compared with two classical elastic buckling solutions and FEM results. Case A: The shell is modeled as a material with an initial modulus ($E=E_{\infty}+E_{1}$). Case B: The shell is modeled with its long-term elastic modulus ($E=E_{\infty}$). The comparison indicates that the buckling pressure of the viscoelastic cylinder is within the expected range between the two elastic cases.
@@
The results presented for both the cantilever beam and the thin cylindrical shell confirm that the proposed PINN framework is a viable and accurate tool for predicting creep buckling phenomena of viscoelastic structures. The findings consistently suggest that PINN can capture these instability-driven deformations naturally through their intrinsic training dynamics, potentially streamlining stability analysis compared to traditional methods requiring explicit perturbations. This capability highlights the significant potential of PINN for analyzing time-dependent instabilities in diverse engineering and biomechanics applications.
@@ 5 PINN for Viscoelastic Growth and Buckling of Soft Biological Tissues
In the implementation of energy-based PINN, which minimizes the system’s potential energy functional by approximating the displacement field, the total potential energy at time $t$ is written as
@@
where the deformation gradient tensor is decomposed using Eq. (15), $I^{a}_{1}=\text{trace}(\mathbf{F}_{a}\cdot\mathbf{F}_{a}^{T})$, $J_{a}=\det(\mathbf{F}_{a})$, and $J_{g}=\mathbf{F}_{g}$. The solutions of the displacement field $u(t)$, stress field $\sigma(t)$ and growth ratio $g_{i}(t)$ are found at discrete time points $t_{k}$. The process repeats for the next time step. To advance the growth ratio $g_{i}$ from time $t$ to $t+\delta t$, it is calculated using Eq. (16), i.e.,
@@ 5.1 Uniform Growth of a Cylindrical Tissue
Figure 7: Cylinder symmetric growth with outer surface fixed. The growth process will proceed to the end, i.e., when the accumulated residual stress is equal to $b^{g}$. (a) Many biological tissues can be seen as a cylinder, e.g., the airway and blood vessel. (b) The schematic diagram and discretization of a semi-cylinder. (c) The total displacement at different times with biochemical stress $b^{g}=0.1E_{0}$. (d) The comparison of growth ratio of the case (c) between PINN and FEM (ABAQUS with CPE4RH elements and user subroutine UMAT) results. (d) The comparison of the growth ratio of the case (c). (e) The total displacement at different times with biochemical stress $b^{g}=0.2E_{0}$. (f) The comparison of the growth ratio of the case (e).
@@
Figs. 7c ($b^{g}=0.1E_{0}$) and 7e ($b^{g}=0.2E_{0}$) depict the displacement magnitude at different times, where $\tau=3$ s is the relaxation time. The color plots confirm that the cylindrical tissue expands uniformly in the radial and axial directions, reflecting isotropic growth. No signs of instability (wrinkling or buckling) are observed, as expected for uniform growth conditions. In Figs. 7d and 7f, we compare the growth ratio over time obtained from the PINN approach against FEM simulations. The results show excellent agreement between PINN and FEM, demonstrating the accuracy and robustness of the proposed approach. For tissues such as blood vessels or airways, uniform growth typically corresponds to healthy, symmetric development, free from pathological remodeling or localized aneurysmal expansions. PINN provide a mesh-free and flexible framework to study these scenarios without requiring heavy computational meshing and complex programming (e.g., UMAT for ABAQUS). Although no instability occurs under uniform growth, the development process may introduce localized deformations, potentially leading to wrinkling or bifurcation. This should consider differential growth in different directions, which can trigger complex pattern formations.

### v2

@@ 3 Numerical Implementation of Viscoelasticity via Energy-based PINN
This section details the numerical framework used to simulate the nonlinear viscoelastic phenomena described in the previous section. We employ an energy-based PINN approach, also known as the Deep Energy Method (DEM) [Samaniego.2020, Abueidda.2022, Zhuang.2021, Bai.2025]. This method seeks the admissible displacement field that minimizes the total potential energy of the system. To handle the inherent time-dependency effect of viscoelasticity and growth, we utilize an incremental, time-stepping procedure, illustrated schematically in Fig. 3.
@@ 3.1 Energy-Based Formulation for Viscoelasticity
where $N(\cdot)$ is the neural network mapping with trainable parameters $\mathbf{\varphi}$ at time step $t_{i}$, and $\mathbf{x}$ is the position of material points. In practice, the neural network architecture, i.e. the number of hidden layers and the number of nodes per layer, is flexible and can be adjusted based on the PDE complexity [Grossmann.2024, Nguyen-Thanh.2020, Li.2024, Bai.2023]. In this study, the neural network has three hidden layers and each layer contains twenty nodes.
@@
using the gradient descent method (e.g., Adam [Kingma.2014]). Once converged, the displacement field $\mathbf{u}(\mathbf{x},t_{i})$ is obtained, and the corresponding stress tensor can be calculated using the constitutive equation. Then, the simulation advances to the next time step $t_{i+1}$ with increment $\Delta t$. This incremental process allows simulation of the full time-dependent viscoelastic response.
@@ 3.2 Implementation of Boundary Conditions
Traction boundary conditions are naturally incorporated through the external work term in the potential energy functional $\Pi$. For displacement (essential) boundary conditions, there are two main approaches to impose within energy-based PINN: the soft and hard ways. The soft way means adding a penalty term to the loss function that penalizes deviations from the prescribed displacement [Barrett.1986, Bai.2025]. We use the hard enforcement to impose the displacement boundary conditions. This method imposes the boundary conditions exactly without adding penalty terms to the loss function, which can be beneficial for training convergence [Bai.2025]. $\mathbf{u}(\mathbf{x},t)=\mathbf{\bar{u}}(\mathbf{x},t)$ for $\mathbf{x}\in\Gamma(\mathbf{x})$ can satisfy by modifying the displacement filed output by PINN through distance functions:
@@
where $\odot$ means the element-wise product, and $b$ is an approximate distance function to the boundary $\Gamma(\mathbf{x})$ which denotes the shortest distance of a point $\mathbf{x}$ to the essential boundary. The distance function is non-negative and $b(\mathbf{x})=0$ only when $\mathbf{x}\in\Gamma(\mathbf{x})$. For the simple geometry with explicit boundaries, the method in Eq. (20) is easy to implement. For complex geometries, the boundary conditions can imposed using distance functions [Sukumar.2022]. Different methods to impose displacement boundary conditions on PINN have been listed and compared in [Berrone.2023].
@@ 3.3 Benchmark Test and Discussion
Figure 4: Viscoelastic creep and stress relaxation cases solved by PINN. (a) The schematic diagram of stress relaxation of a cantilever. The left boundary is fixed, and the right is fixed horizontally after stretching for a displacement $\Delta L$. (b) Comparison of the stress relaxation results at the time $t=3\tau$ from FEM and PINN. (c) The schematic diagram of creep of a cantilever. The left boundary is fixed, and the right boundary is subjected to a horizontal uniformly distributed tensile force. (d) Comparison of the creep results at the time $t=3\tau$ from FEM and PINN. (e) The training dynamics of the energy-based PINN. The potential energy loss decreases as training progress. Small oscillations may occur during the convergence process, which are brought in by the optimization algorithm itself.
@@
The viscoelastic creep displacement obtained from the PINN agree well with the finite element results (Fig. 4d). These benchmark results demonstrate that the implemented energy-based PINN framework, using an incremental time-stepping approach, accurately captures the fundamental time-dependent responses of viscoelastic materials.
@@ 3.4 Oscillatory Convergence of Neural Network
An interesting characteristic observed during the training of energy-based PINN is the oscillatory behavior in the convergence of the loss function or predicted displacements (Fig. 4e). These oscillations arise from the dynamics of the gradient descendant optimizer (e.g., SGD and Adam), which is very common during the deep learning. Such oscillations can facilitate escape from poor local minima but they may impede stable convergence during later optimization stages. In contrast, such oscillatory behavior may be advantageous for capturing phenomena like buckling in viscoelastic materials.
Classical numerical methods often require explicit artificial perturbations or initial imperfections to initiate buckling bifurcation from a perfect equilibrium path. We hypothesize that the inherent training oscillations within the energy-based PINN can act as a form of natural, internal perturbation. As the optimizer explores the energy landscape near a bifurcation point (where the landscape topology changes), these numerical oscillations can be sufficient to nudge the solution off the unstable primary path and guide it towards the physically stable, lower-energy buckled configuration. The potential applications are explored further in next sections.
@@ 4.1 Creep Buckling of a Viscoelastic Cantilever Beam
Figure 5: Creep buckling of a cantilever beam under axial compression. (a) Schematic diagram of a compressed cantilever beam. (b) Uniform discretization of the beam. (c) Displacement in the y-direction of the creep buckling cantilever beam at different times obtained from PINN under pressure $2\text{e}-2$ Pa, with Adam optimizer and learning rate $5\text{e}-3$. (d) Displacement in y-direction obtained from the Fourier feature-PINN ($F$-PINN). The FF-PINN has three layers, and the feature number of each layer is 64 in this study. (e) Deformed cantilever beam obtained from FEM under pressure $2\text{e}-2$ Pa, with initial geometric imperfections. (f) Pointwise error of the y-displacement obtained from PINN and FEM at $t=2\tau$. The error corresponds to the square error. (g) A comparison of the time-deflection curves obtained from PINN and FEM.
@@
The evolution of the deformation over time is shown in Fig. 5c-e. As creep deformation progresses, the beam undergoes significant lateral displacement, leading to buckling. To assess the effect of Fourier feature embeddings on representational capacity, we computed the solution using a Fourier-Feature PINN (FF-PINN) [WANG2021113938], as shown in Fig. 5d. The FF-PINN produces deformation fields that are qualitatively similar to those of the baseline PINN and does not lead to a noticeable improvement for this problem. This is expected because the physical response is dominated by low-order global buckling modes, so the added high-frequency expressivity has limited influence on accuracy. The creep buckling with FEM is predicted by adding some initial geometric imperfections (Fig. 5e) under the same axial pressure. A quantitative comparison of the methods is provided in Fig. 5f, which shows the squared pointwise error fields between the PINN prediction and FEM solution at $t=2\tau$. Fig. 5g compares the normalized mid-span deflection over time. The time–deflection curves of the PINN and FEM do not coincide at every time point once the structure enters the post-buckling regime. This discrepancy is expected: after the onset of buckling, the mechanical system becomes highly sensitive to small perturbations, and the post-buckling path is no longer unique. In FEM, the specific imperfection introduced to trigger buckling determines which post-buckling branch is followed. While in the PINN approach, the optimizer-induced perturbations naturally select a different, though mechanically admissible, buckling trajectory. Because these post-critical paths correspond to distinct but energetically similar equilibrium branches, perfect time-point agreement between PINN and FEM is not expected. Importantly, both methods predict the same buckling onset, overall mode shape, and qualitative evolution, which is a meaningful benchmark for creep-induced instability.
The PINN solution captures the gradual accumulation of strain and the onset of instability. To validate the predicted buckling behavior, we compare the eigenvalue of the viscoelastic cantilever with two reference cases obtained from FEM simulations (ABAQUS, plane strain analysis with CPE8R elements): one is modeled as an elastic material with initial modulus $E_{0}=E_{\infty}+E_{1}$, representing the upper bound of stiffness, and another is modeled with its long-term modulus $E=E_{\infty}$, serving as a baseline reference. The eigenvalue of a beam with the infinity modulus ($E_{\infty}$) is obtained from finite element simulation, which is $9.36\text{e}-3$, and the beam with initial stiffness ($E_{0}$) is $2.34\text{e}-2$. The buckling pressure $2\text{e}-2$ for viscoelastic cantilever is between these two elastic conditions. These predictions confirm that the buckling pressure of the viscoelastic cantilever falls within the expected range between two elastic cases. Crucially, the buckling behavior in the PINN simulation emerged without introducing any explicit geometric imperfections or transient load perturbations. The transition to the buckled state appears to be triggered by the inherent dynamics of the optimizer exploring the solution space near the bifurcation point, consistent with the hypothesis outlined in Section 3.4. As supporting evidence, simulations performed with a sub-critical load ($p=5\text{e}-3<p_{\text{cr}}$) showed only minor oscillations or shaking during training without leading to buckling, as detailed in Appendix B. This highlights a potential advantage of PINN for stability analysis.
@@ 4.2 Creep Buckling of a Thin Cylindrical Shell under External Pressure
The discretization of the shell domain is shown in Fig. 6b, where the uniform sampling points are used to resolve the spatial dependence of the displacement field. Sample points on the outer surface of the shell are used to calculate the outer surface displacement. The implement of this problem in PINN is under the cylindrical coordinate. The boundary condition is introduced to the neural network as
@@
Figure 6: Creep buckling of a thin cylinder with outside radial pressure. (a) The schematic diagram of a compressed quarter cantilever cylinder, with the pressure equals to $1\text{e}-2$. (b) The discretization of the cylinder, and the integration at the outer surface. (c) The outer surfaces at different times. (c) Deformed quarter cylinder at different times. The forth order eigenvalue of a beam with the infinity modulus ($E_{\infty}$) is obtained from finite element simulation (Plain strain, CPE8R), which is $6.44\text{e}-3$, and the cylindrical shell with initial stiffness ($E_{0}$) is $1.61\text{e}-2$. The buckling pressure $1\text{e}-2$ for viscoelastic cantilever is between these two elastic conditions.
@@
The evolution of the shell’s deformation is presented in Figs. 6c and 6d. In Fig. 6c, the deformation profile of the outer surface is plotted at different time instances, illustrating the progressive lateral deflection and the onset of buckling. The initially circular shell gradually deviates from its original shape due to viscoelastic creep, eventually leading to instability. The corresponding displacement fields at specific time snapshots are visualized in Fig. 6d. At $t=0^{+}$, the shell exhibits a nearly uniform radial contraction due to the applied pressure. As creep deformation progresses, localized bending deformations emerge, which intensify over time, culminating in a pronounced buckled shape at $t=4\tau$. To validate the PINN predictions, the buckling pressure obtained from the viscoelastic simulation is still compared with two classical elastic buckling solutions and FEM results. Case A: The shell is modeled as an material with initial modulus ($E=E_{\infty}+E_{1}$). Case B: The shell is modeled with its long-term elastic modulus ($E=E_{\infty}$). The comparison indicates that the buckling pressure of the viscoelastic cylinder within the expected range between two elastic cases.
@@
The results presented for both the cantilever beam and the thin cylindrical shell confirm that the proposed PINN framework is a viable and accurate tool for predicting creep buckling phenomena of viscoelastic structures. The findings consistently suggest that PINN can capture these instability-driven deformations naturally through their intrinsic training dynamics, potentially streamlining stability analysis compared to traditional methods requiring explicit perturbations. Furthermore, we also proved theoretically how the Adam optimizer brings in perturabation for buckling (Appendix C). This capability highlights the significant potential of PINN for analyzing time-dependent instabilities in diverse engineering and biomechanics applications.
To evaluate the practical viability of the proposed method, we compare the wall-clock time of the PINN simulations with those of equivalent FEM simulations performed on the same workstation. Table 1 summarizes the computational cost for all benchmark problems. All problems are modelled on a CPU AMD Ryzen 7 5700G (3.80 GHz). Although the current PINN implementation is not computationally faster than FEM, it provides capabilities, such as mesh-free deformation, natural instability capture, and robust handling of evolving geometry, that are difficult for FEM to replicate. These complementary strengths justify PINN development despite the higher computational expense. Additionally, inference of the PINN is extremely fast, requiring only a forward evaluation of the neural network. The trained model can also be fine-tuned for nearby parameter regimes at a very low additional cost, amortizing the initial training effort across multiple related problems.
Table 1: A computational cost comparison of PINN and FEM for viscoelastic creep-buckling
Problem Method Mesh/Network size CPU time for each time step
Cantilever beam buckling FEM 1000 elements < 1 s
[…]
Cylindrical shell buckling FEM 1500 elements < 1 s
[…]
@@ 5 PINN for Viscoelastic Growth and Buckling of Soft Biological Tissues
In the implementation of energy-based PINN which minimizes the system’s potential energy functional by approximating the displacement field, the total potential energy at time $t$ is written as
@@
where the deformation gradient tensor is decomposed using Eq. (15), $I^{a}_{1}=\text{trace}(\mathbf{F}_{a}\cdot\mathbf{F}_{a}^{T})$, $J_{a}=\det(\mathbf{F}_{a})$, and $J_{g}=\mathbf{F}_{g}$. The solutions of displacement field $u(t)$, stress field $\sigma(t)$ and growth ratio $g_{i}(t)$ are found at discrete time points $t_{k}$. The process repeats for the next time step. To advance the growth ratio $g_{i}$ from time $t$ to $t+\delta t$, it is calculated using Eq. (16), i.e.,
@@ 5.1 Uniform Growth of a Cylindrical Tissue
Figure 7: Cylinder symmetric growth with outer surface fixed. The growth process will proceed to the end, i.e., when the accumulated residual stress is equal to $b^{g}$. (a) Many biological tissues can be seen as a cylinder, e.g., the airway and blood vessel. (b) The schematic diagram and discretization of a semi-cylinder. (c) The total displacement at different times with biochemical stress $b^{g}=0.1E_{0}$. (d) The comparison of growth ratio of the case (c) between PINN and FEM (ABAQUS with CPE4RH elements and user subroutine UMAT) results. (d) The comparison of growth ratio of the case (c). (e) The total displacement at different times with biochemical stress $b^{g}=0.2E_{0}$. (f) The comparison of growth ratio of the case (e).
@@
Figs. 7c ($b^{g}=0.1E_{0}$) and 7e ($b^{g}=0.2E_{0}$) depict the displacement magnitude at different times, where $\tau=3$ s is the relaxation time. The color plots confirm that the cylindrical tissue expands uniformly in the radial and axial directions, reflecting isotropic growth. No signs of instability (wrinkling or buckling) are observed, as expected for uniform growth conditions. In Figs. 7d and 7f, we compare the growth ratio over time obtained from the PINN approach against FEM simulations. The results show excellent agreement between PINN and FEM, demonstrating the accuracy and robustness of the proposed approach. For tissues such as blood vessels or airways, uniform growth typically corresponds to healthy, symmetric development, free from pathological remodeling or localized aneurysmal expansions. PINN provide a mesh-free and flexible framework to study these scenarios without requiring heavy computational meshing and complex programming (e.g., UMAT for ABAQUS). Although no instability occurs under uniform growth, the development process may introduce localized deformations, potentially leading to wrinkling or bifurcation. This should consider differential growth at different directions, which can trigger complex pattern formations.